OpenStudy (anonymous):

12(x-y)^-1/2 ------------ = 2(x-y)^1/2

OpenStudy (anonymous):

6/(x-y)

OpenStudy (zzr0ck3r):

(12/2)(x-y)^(-1/2-1/2) = 6(x-y)^(-1) = 6/(x-y)

OpenStudy (anonymous):

Thanks!

OpenStudy (zzr0ck3r):

np